Five times the measure of an angle is twelve less than its supplement. Find the measure of an angle

Answer by rm29924(97)

Let x=the measure of an angle
break the word problem up
five times the measure of an angle would be 5x
five times the measure of an angle 5x = twelve less than it's supplement
an angle's supplement added to the angle = 180 degrees
so the angle's supplement is 180-x
5x=180-x-12
5x=168-x
6x=168
x=28
check answer:
5(28)=180-28-12
140=140